OpenStudy (anonymous):

4x - 6y = 18, find y when x = 3 my answer is -1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

lets check \[4\times 3-6y=18\\ 12-6y=18\\ -6y=6\\ y=-1\]yup

OpenStudy (anonymous):